For-Profit
20% of marketing expenses are fixed; the rest are variable; all other expenses are fixed
Next year they plan to double their fixed marketing expenses and as a result will increase ticket sales by 25%.
Gift shop sales vary directly with ticket sales; gift shop COGS is 50% of sales but is expected to grow to 55% of sales.
For-Profit Dinosour Museum
Income Statement
This Yr Next Yr Budget Reason
Ticket sales $600,000
Gift shop sales 200,000
Total sales 800,000
Salaries & benefits 100,000
Gift shop COGS 100,000
Marketing 100,000
Maintenance 90,000
Skeleton depreciation 20,000
Building lease 20,000
Utilities 40,000
Interest on debt 40,000
Total expense 510,000
Before tax income 290,000
Tax @ 40% (116,000) (165,000)
After-tax income 174,000 247,500
  • complete next year's budgeted figures for each entity within the same Excel worksheet (yellow cells)
  • include your reasoning for most of the budgeted figures (green cells)

Because of a bad recession, government grants were not cut by 50% but are eliminated completely for next year's budget and fundraising efforts—despite an increase in expenses—did not improve at all.
Answer #1
For-Profit Dinosour Museum
Income Statement
This Yr Next Yr Budget Reason
Ticket sales 600000 750000 Increase by 25% due to increased fixed marketing expenses
Gift shop sales 200000 250000 Gift shop sales directly vary with ticket sales. Gift shop sales = (200000*(750000/600000))
Total sales 800000 1000000
Salaries & benefits 100000 100000 Nothing has been given about increase in this expenditure
Gift shop COGS 100000 137500 Gift shop COGS is going to be 55% of sales i.e. 55% of 250000
Marketing 100000 140000 20% of marketing expenses are fixed. Fixed marketing expenses are doubled now. Fixed marketing expneses = 100000*20%*2 =40000
Variable marketing expenses = (100000-20000)*(1000000/800000) = 100000
Total expenses = 40000+100000=140000
Maintenance 90000 90000 Nothing has been given about increase in this expenditure. As they are fixed, they remain the same
Skeleton depreciation 20000 20000 Nothing has been given about increase in this expenditure. As they are fixed, they remain the same
Building lease 20000 20000 Nothing has been given about increase in this expenditure. As they are fixed, they remain the same
Utilities 40000 40000 Nothing has been given about increase in this expenditure. As they are fixed, they remain the same
Interest on debt 40000 40000 Nothing has been given about increase in this expenditure. As they are fixed, they remain the same
Total expense 510000 587500
Before tax income 290000 412500
Tax @ 40% -116000 165000
After-tax income 174000 247500